OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. Write your answer in set brackets with no spaces, such as {-1,2}. x2 - 4x - 32 = 0

OpenStudy (anonymous):

x^2-4x+4=32+4 (x-2)(x-2)=36 (x-2)^2=36 x-2= +/- 6 x=8,-4 {-4,8}